# yijia#1 [Yijia: #1] Large Language Models Can Be Good Privacy Protection Learners

### v1

@@ 5.3 Evaluation Metrics
We use the Q/A task as the validation protocol. To validate how good the domain knowledge in the fine-tuning texts is effectively incorporated into the resulting language model, we adopt the popularly used ROUGE-1, ROUGE-2, ROUGE-L (Lin, 2004) and BERTScore (Zhang* et al., 2020) to evaluate the answer quality in the testing phase. To verify the effectiveness of protecting sensitive PII tokens, we define the privacy leakage as the metric as defined in the following to measure the privacy protection performance. The detailed definition is also included in the Appendix A.2.
@@
5.4 Description of different methods validated
To demonstrate the efficiency of our methods, we compare the proposed strategies. Besides, we also provide an additional approach as our baseline. Since prepending instructions ahead of the model’s input can tune the model to follow instructions (Wang et al., 2022; Taori et al., 2023; Han et al., 2023), we define the Vanilla tuning (visualized in Appendix B) borrowing this idea as our baseline. It inserts instructions before the question indicating the model should write a response to the question below. Removal denotes the strategy of extracting PIIs from the corpus. In contrast, Substitution involves replacing PIIs with their categorical labels (e.g. NAME, ORGANIZATION, URL, ADDRESS). Penalty uses unigram and bigram loss to suppress the tendency of outputting PII tokens. Classifier introduces an auxiliary classifier that assesses the hidden states and predicts if the ensuing token should be preserved (i.e., not displayed in the generated text). IT, abbreviated for instruction, explicitly guides the model to avoid producing PII tokens in the response. Both and refer to instruction tuning with specific (positive/negative) cases: PN pertains to the positive-negative case order, and NP to the negative-positive case order. The “Instruction with Case” chart in Figure 1 showcases , while for , the cases are inverted. Furthermore, the subscripts 1/2 in delineate different instructions (refer to Appendix C.4 for details).
@@ 5.5 Results and Analysis
In this experimental analysis, we assess the performance of different methods for enhancing privacy in language models while considering their impact on knowledge retention as measured by ROUGE scores and BERTScore (). From Table 1, 2 and 3, we can observe that the score of the Vanilla method is high, which suggests that it is susceptible to privacy breaches due to its reliance on the entire training textual data without any privacy protection strategy. The “Removal” and “Substitution” methods effectively safeguard privacy. They both focus on privacy protection by actively removing sensitive information from the model’s knowledge base. The removal of sensitive information significantly reduces the knowledge retained by the model. The and ROUGE scores are observed a substantial drop due to the removal of data, resulting in reduced language understanding and generation abilities. We also note that the penalty-based approach can effectively safeguard privacy. However, since it imposes constraints on selectively forgetting previous knowledge, it may inadvertently alter existing knowledge and encourage the model to generate alternative tokens for PIIs, potentially distorting the original information. As a result, its performance in injecting knowledge experiences a slight reduction in some datasets. The “Classifier” approach yields moderate results in terms of privacy protection, highlighting the inherent challenge in training a contextual classifier of this nature. The DPO approach first adopts the Vanilla tuning (SFT) without privacy protection, then further fine-tune the model to instruct the model in concealing sensitive PII without the requirement of learning a reward model. From the results, it is evident that DPO can enhance privacy protection through preference-based tuning. However, its performance is suboptimal, possibly because it often requires a larger dataset of user preferences for effective training, and the potential issue of reward hacking. Our experiments indicate that the instruction tuning with examples approach, which utilizes instructions and examples to fine-tune the model, strikes a favorable balance between performance and objectives such as privacy protection, user-defined information preservation, and alignment with human preferences. The results indicate that letting the model “see” the preferred example and undesired examples is helpful in model alignment. Such an approach enables the model to capture the information from the corpus and learn what should not be released. This underscores the potential for LLMs as privacy protection learners.
@@
5.6 Curve of Knowledge Injection and PII leakage vs. learning process
In this section, we conduct an analysis of the ROUGE/BERTScore and the Privacy Leakage Score in relation to the training steps. Our aim is to assess whether our two primary learning objectives are effectively achieved throughout the training process. Initially, in Figure. 2 that visualizes the training of (more results are included in the Appendix), we observe that as the LM undergoes the training process, we witness a notable trend: the injection of knowledge into the LM steadily increases. This infusion of knowledge corresponds to a progressive rise in both ROUGE and BERTScore, ultimately leading to a stabilization, or convergence, of these metrics. Simultaneously, the Privacy Leakage Score exhibits an intriguing behavior. At the outset of the learning process, it experiences an upward trajectory. This ascent is a direct consequence of the LM ingesting more knowledge, including private tokens, inadvertently learning about sensitive information. However, as training continues, a pivotal shift occurs. The LM’s instruction to conceal privacy-related information gradually takes effect, resulting in a discernible decrease in the Privacy Leakage Score. In summary, Figure. 2 offers a compelling visualization of the evolving relationship between knowledge injection, linguistic performance (ROUGE/BERTScore), and privacy protection () as the LM matures throughout its training steps. It underscores the dynamic equilibrium between knowledge acquisition and safeguarding sensitive data, emphasizing the importance of a well-orchestrated learning process to achieve both objectives.
@@ 6 Conclusion
In this paper, we present a comprehensive exploration of strategies for fine-tuning Large Language Models (LLMs) to incorporate domain-specific knowledge while upholding data privacy, particularly in safeguarding sensitive Personally Identifiable Information (PII). We introduced the novel concept of Privacy Protection Language Models (PPLMs) and provided a theoretical analysis to guide model design. Our extensive experiments underscore the effectiveness of our approach, with instruction-based tuning emerging as a promising method to simultaneously protect private data and enhance the model’s knowledge. This study highlights the potential for LLMs to serve as adept privacy protection learners, bridging the gap between domain-specific expertise and data privacy. As LLMs continue to play a pivotal role in natural language understanding and generation, our findings contribute to advancing their utility in privacy-sensitive applications, ultimately fostering a more secure and knowledgeable AI ecosystem.
Ethics Statement
All contributing authors of this paper confirm that they have read and pledged to uphold the ICLR Code of Ethics.
Reproducibility Statement
All specifics regarding the datasets and our experimental configurations can be found in Appendices A.1 and C. The source code and scripts for experiments, available in an anonymized form, can be accessed at https://anonymous.4open.science/r/PPLM/.
@@ A.2 Metrics
Appendix B Illustration of Vanilla Tuning and Corpus Curation
@@
Appendix C Experiment Details.
C.1 Hardware and Implementations
In this paper, we implemented our method on two Linux servers with 4 NVIDIA A100 GPUs, each with 80GB of memory. The CUDA version is 12.2 and the Driver version is 535.54.03. We used Python 3.10.12 and Pytorch 2.0.1 (Paszke et al., 2019) to construct our project.
C.2 Dataset and Hyperparameters
@@
C.3 Utility and Privacy Curve
@@
To compare vanilla tuning with instruction tuning using positive-negative cases (), we plotted utility metrics (ROUGE/BERTScore) and against the number of training steps (as shown in Figure. 4 and Figure. 2). With vanilla tuning (Figure. 2), as training progresses, the LLM’s performance improves. However, it is accompanied by an increase in privacy leakage. Such a trend corroborates our intuition that, as the LLM assimilates information, it also inadvertently memorizes PII tokens from the corpus. When it comes to instruction tuning with positive-negative cases (Figure. 2), the utility curve exhibits a trajectory akin to vanilla tuning. However, privacy leakage increases initially but eventually declines. This suggests that, by employing instruction combined with positive-negative cases, LLMs can be trained to be good privacy learners.
C.4 Instruction Templates
@@
In this section, we list the detailed two instruction tuning templates as shown in Table. 6.
@@
Appendix D Case Study
@@
In this section, we present case studies featuring exemplary responses generated from the Wikidoc Patient Information Dataset. We compare responses generated by the vanilla model (original) with those generated by a model equipped with a token classifier and responses generated by the vanilla model (original) versus the model trained using the instruction tuning (IT) strategy. By examining the results in Tables 7 and 8, we can discern the efficacy of the proposed methodologies in safeguarding privacy.

### v2

@@ 5.3 Evaluation Metrics
We use the Q/A task as the validation protocol. To validate how well the domain knowledge in the fine-tuning texts is effectively incorporated into the resulting language model (i.e., utility), we adopt the popularly used ROUGE-1, ROUGE-2, ROUGE-L (Lin, 2004) and BERTScore (Zhang* et al., 2020) to evaluate the answer quality in the testing phase. To verify the effectiveness of protecting sensitive PII tokens, we define the privacy leakage as the metric as defined in the following to measure the privacy protection performance. The detailed definition is also included in the Appendix A.2.
@@
5.4 Different Methods Validated
To demonstrate the efficiency of our methods, we compare the proposed strategies. Besides, we also provide an additional approach as our baseline. Since prepending instructions ahead of the model’s input can tune the model to follow instructions (Wang et al., 2022; Taori et al., 2023; Han et al., 2023), we define the Vanilla tuning (visualized in Appendix C) borrowing this idea as our baseline. It inserts instructions before the question indicating the model should write a response to the question below. Removal denotes the strategy of extracting PIIs from the corpus. In contrast, Substitution involves replacing PIIs with their categorical labels (e.g. NAME, ORGANIZATION, URL, ADDRESS). Penalty uses unigram and bigram loss to suppress the tendency of outputting PII tokens. Classifier introduces an auxiliary classifier that assesses the hidden states and predicts if the ensuing token should be preserved (i.e., not displayed in the generated text). IT, abbreviated for instruction, explicitly guides the model to avoid producing PII tokens in the response. Both and refer to instruction tuning with specific (positive/negative) cases: PN pertains to the positive-negative case order, and NP to the negative-positive case order. The “Instruction with Cas” chart in Figure 1 showcases , while for , the cases are inverted. Furthermore, the subscripts 1/2 in delineate different instructions (Appendix D.6).
@@ 5.5 Results and Analysis
In this experimental analysis, we assess the performance of different methods for enhancing privacy in language models while considering their impact on knowledge retention as measured by ROUGE scores and BERTScore (). In Appendix D.5, we analyze the ROUGE, BERTScore, and Privacy Leakage Score concerning the training steps to assess whether our two learning objectives are effectively achieved throughout the training process.
@@
In Tables 1, 2, and 3, the high score for the Vanilla method indicates its vulnerability to privacy breaches, as it uses all training text data without privacy preservation. The “Removal” and “Substitution” methods effectively safeguard privacy. They both focus on privacy protection by actively removing sensitive information from the model’s knowledge base. The removal of sensitive information significantly reduces the knowledge retained by the model. The and ROUGE scores are observed to suffer a substantial drop due to the removal of data, resulting in reduced language understanding and generation abilities. We also note that the penalty-based approach can effectively safeguard privacy.
@@
| Strategy | ROUGE-1 | ROUGE-2 | ROUGE-L | BERTScore | ||||||||
| Vanilla | 0.637 | 0.5743 | 0.6235 | 0.8699 | 0.0778 | - | 0.0752 | - | 0.0782 | - | 0.0724 | - |
| Removal | 0.6148 | 0.5575 | 0.6115 | 0.8390 | 0.0410 | -47.30 | 0.0394 | -47.61 | 0.0423 | -45.91 | 0.0419 | -42.13 |
| Substitution | 0.6291 | 0.5234 | 0.6217 | 0.8576 | 0.0420 | -46.02 | 0.0418 | -44.41 | 0.0446 | -42.97 | 0.0419 | -42.13 |
| IT | 0.6395 | 0.5429 | 0.6253 | 0.8686 | 0.0449 | -42.29 | 0.0418 | -44.41 | 0.0449 | -42.58 | 0.0421 | -41.85 |
| 0.6497 | 0.5591 | 0.6346 | 0.8696 | 0.0395 | -49.23 | 0.0397 | -47.21 | 0.0419 | -46.42 | 0.0411 | -43.23 | |
| 0.6324 | 0.5569 | 0.6222 | 0.869 | 0.0404 | -48.07 | 0.0403 | -46.41 | 0.0421 | -46.16 | 0.0413 | -42.96 | |
| 0.6321 | 0.5740 | 0.6234 | 0.8605 | 0.0411 | -47.17 | 0.0412 | -45.21 | 0.0431 | -44.88 | 0.0414 | -42.82 | |
| 0.6335 | 0.5761 | 0.6201 | 0.8657 | 0.0406 | -47.81 | 0.0408 | -45.74 | 0.0412 | -47.31 | 0.0416 | -42.54 |
Selective forgetting constraints in models may inadvertently alter existing knowledge, leading to token alterations for PIIs and possibly distorting original information, slightly reducing performance in some datasets. The “Classifier” approach offers moderate privacy protection results, reflecting the challenge in training contextual classifiers. DPO starts with Vanilla tuning (SFT) without privacy measures, then fine-tunes for PII concealment without a reward model. While DPO boosts privacy through preference-based tuning, its effectiveness is limited, often needing a larger dataset of user preferences and facing reward hacking issues.
Experiments show that instruction tuning with examples, using instructions and examples for fine-tuning, achieves a good balance between performance, privacy, information preservation, and alignment with human preferences. This method, letting the model “see” and “learn” from both preferred and undesired examples, helps in aligning the model. It enables the model to understand what information to withhold, highlighting the potential of LLMs in privacy protection learning.
We also plot the Pareto frontier in Figure 2(a) and 2(b) to evaluate both utility and privacy preservation on medical_flashcards dataset for LLaMA2-7B and LLaMA2-13B, respectively. More results are reported in Appendix D.4. It is evident that the instruction-based approaches consistently align with the Pareto frontier (IT methods constitute the border of the frontier). Such a phenomenon indicates that employing instructions supplemented by both positive and negative examples achieves the optimal trade-off between performance (utility) and privacy protection of PIIs.
5.6 Performance on Different Types of PIIs
To validate the performance of our approaches on different types of PIIs, we have conducted further experiments on the newly synthesized dataset. The dataset, named Privacy QA (PQA) Dataset, was synthesized using GPT-4. The PQA dataset contains a wider range of entities, including Names, Emails, Addresses, and SSNs. PQA is accessible at the anonymous link333https://github.com/Yijia-Xiao/PPLM/ft_datasets/data/PQA.csv. The categorization helps assess the protection effectiveness for each PII type. For instance, SSN leaks are generally more critical than name leaks. We performed experiments on the Privacy QA dataset, evaluating the protection ratios across these PII categories respectively. The evaluation is performed on LLaMA2-7B and results are provided in Table 4. The results show that the instruction tuning approaches can well protect different types of PIIs while providing good knowledge injections.
@@ 6 Conclusion
In this paper, we present a comprehensive exploration of strategies for fine-tuning Large Language Models (LLMs) to incorporate domain-specific knowledge while upholding data privacy, particularly in safeguarding sensitive Personally Identifiable Information (PII). We introduced the novel concept of Contextual Privacy Protection Language Models (CPPLMs) and provided a theoretical analysis to guide model design. Our extensive experiments underscore the effectiveness of our approach, with instruction-based tuning emerging as a promising method to simultaneously protect private data and enhance the model’s knowledge. This study highlights the potential for LLMs to serve as adept privacy protection learners, bridging the gap between domain-specific expertise and data privacy. As LLMs continue to play a pivotal role in natural language understanding and generation, our findings contribute to advancing their utility in privacy-sensitive applications.
Limitations
CPPLM explores privacy preservation in large language models. It is important to note that in our dataset, personally identifiable information (PII) is identified using the scrubadub toolkit. Such a tagging process may not fully represent real-world deployment scenarios, where users can customize privacy preferences. Companies and data owners can employ the CPPLM pipeline to teach language models contextual privacy from annotated positive-negative pairs. Since there is no universal rule for detecting PIIs, privacy definitions vary across scenarios. Therefore, our focus is on demonstrating the language model’s ability to learn contextual PII. For instance, a clinical company wanting to protect specific PIIs can annotate datasets and follow our proposed method. Even end-users may define what PII means in their data’s context during language model tuning or training. In summary, the CPPLM pipeline is versatile and adaptable to various privacy-related scenarios and tasks, such as detoxifying language models.
Acknowledgments
This work was supported by NSF grants #2200274, #2106859, #2312501, NIH grants #U54HG012517, #U24DK097771.
Appendix: Large Language Models Can Be Contextual Privacy Protection Learners
@@ A.2 Metrics
Appendix B Additional Related Work
Pretraining with Preferences. Another solution is to maintain the content, but use redesigned loss/conditional tags to control the information injected into the LLMs. Pretraining with conditional human preference scores can offer a Pareto-optimal and simple approach to reduce the undesirable content by up to an order of magnitude. Korbak et al. (2023) compared with the classical pretraining approach. While pretraining LLMs conditioned under annotation scores can offer better performance in the human preferences aspect. Since human preferences are injected into the models during the pretraining stage, the models are biased toward those preferences once they are trained. With the expanding size of LLMs, they become increasingly resistant to forgetting their training data (Carlini et al., 2022; Vu et al., 2022; Ramasesh et al., 2022; Korbak et al., 2023). In other words, pretraining large language models conditioned under preference score sacrifices some flexibility. Still, it is undeniable that it can provide much better alignment with human preferences compared with the classical pretraining schema.
Appendix C Illustration of Vanilla Tuning and Corpus Curation
@@
Appendix D Experiment Details.
D.1 Hardware and Implementations
In this paper, we implemented our method on two Linux servers with 4 NVIDIA A100 GPUs, each with 80GB of memory. The CUDA version is 12.2 and the Driver version is 535.54.03. We used Python 3.10.12 and Pytorch 2.0.1 (Paszke et al., 2019) to construct our project. The fine-tuning of LLaMA models takes 20 hours on average.
D.2 Dataset and Hyperparameters
@@
D.3 Comparison with Other Methods
In this section, we compare our proposed methods against several privacy-preserving techniques, including Vanilla, Removal, Substitution, and Private Transformer, as well as our IT_PN and IT_NP strategies. The comparison is made on the PQA dataset in terms of both performance (ROUGE and BERTScore metrics) and privacy protection across multiple Personally Identifiable Information (PII) categories such as Name, Email, Address, and SSN.
As shown in Table 7, our methods demonstrate competitive performance in terms of utility, while also providing notable improvements in privacy protection. The IT_PN method, in particular, consistently outperforms the baseline strategies in terms of reducing privacy leakage, achieving a reduction of up to 49.26% in the S_priv:Name metric compared to the Vanilla model. Furthermore, IT_PN maintains a high level of utility, reflected in its ROUGE and BERTScore values, which are comparable to or better than those of the Private Transformer approach.
| Strategy | ROUGE-1 | ROUGE-2 | ROUGE-L | BERTScore | ||||||||
| Vanilla | 0.3342 | 0.2174 | 0.3297 | 0.8162 | 0.1082 | - | 0.1024 | - | 0.1103 | - | 0.0992 | - |
| Removal | 0.2947 | 0.2071 | 0.3092 | 0.7983 | 0.0558 | -48.43% | 0.0536 | -47.66% | 0.0573 | -48.05% | 0.0568 | -42.75% |
| Substitution | 0.2983 | 0.2073 | 0.3173 | 0.8012 | 0.0572 | -47.13% | 0.0569 | -44.43% | 0.0586 | -46.87% | 0.0568 | -42.75% |
| Private Transformer | 0.3172 | 0.2096 | 0.3192 | 0.8119 | 0.0551 | -49.08% | 0.0554 | -45.90% | 0.0572 | -48.14% | 0.0569 | -42.65% |
| IT_PN | 0.3273 | 0.2112 | 0.3221 | 0.8101 | 0.0549 | -49.26% | 0.0551 | -46.19% | 0.0570 | -48.32% | 0.0570 | -42.55% |
| IT_NP | 0.3261 | 0.2162 | 0.3252 | 0.8132 | 0.0563 | -47.97% | 0.0561 | -45.21% | 0.0575 | -47.87% | 0.0573 | -42.24% |
D.4 Pareto Frontier of Utility and Privacy Protection
We also report the pareto frontier of Utility and Privacy Protection in Figure 4, 5, 6, 7, 8 and 8, respectively, to evaluate both performance and privacy preservation. It is obvious that the instruction-based approaches consistently align with the Pareto frontier (IT methods constitute the border of the frontier). Such a phenomenon indicates that employing instructions supplemented by both positive and negative examples achieves the optimal trade-off between performance (utility) and privacy protection of PIIs. The outcomes strongly support our position that LLMs can be good contextual privacy protection learners.
D.5 Curve of Knowledge Injection and PII Leakage vs. Learning Process
In this section, we analyze the ROUGE, BERTScore, and Privacy Leakage Score concerning the training steps. We aim to assess whether our two primary learning objectives are effectively achieved throughout the training process. Initially, in Figure. 10 that visualizes the training of , we observe that as the LM undergoes the training process, we witness a notable trend: the injection of knowledge into the LM steadily increases. This infusion of knowledge corresponds to a progressive rise in both ROUGE and BERTScore, ultimately leading to a stabilization, or convergence, of these metrics. Simultaneously, the Privacy Leakage Score exhibits an intriguing behavior. At the outset of the learning process, it experiences an upward trajectory. This ascent is a direct consequence of the LM ingesting more knowledge, including private tokens, inadvertently learning about sensitive information. However, as training continues, a pivotal shift occurs. The LM’s instruction to conceal privacy-related information gradually takes effect, resulting in a discernible decrease in the Privacy Leakage Score. In summary, Figure. 10 offers a compelling visualization of the evolving relationship between knowledge injection, linguistic performance (ROUGE/BERTScore), and privacy protection () as the LM matures throughout its training steps. It underscores the dynamic equilibrium between knowledge acquisition and safeguarding sensitive data, emphasizing the importance of a well-orchestrated learning process to achieve both objectives.
@@ D.4 Pareto Frontier of Utility and Privacy Protection
To compare vanilla tuning with instruction tuning using positive-negative cases (), we plotted utility metrics (ROUGE/BERTScore) and against the number of training steps (as shown in Figure. 11). With vanilla tuning, as training progresses, the LLM’s performance improves. However, it is accompanied by an increase in privacy leakage. Such a trend corroborates our intuition that, as the LLM assimilates information, it also inadvertently memorizes PII tokens from the corpus. When it comes to instruction tuning with positive-negative cases (Figure. 10), the utility curve exhibits a trajectory akin to vanilla tuning. However, privacy leakage increases initially but eventually declines. This suggests that, by employing instruction combined with positive-negative cases, LLMs can be trained to be good contextual privacy learners.
D.6 Instruction Templates
@@
In this section, we list the detailed two instruction tuning templates as shown in Table. 8 and Table. 9.
@@
Appendix E Case Study
@@
In this section, we present case studies featuring exemplary responses generated from the Wikidoc Patient Information Dataset. We compare responses generated by the vanilla model (original) with those generated by a model equipped with a token classifier and responses generated by the vanilla model (original) versus the model trained using the instruction tuning (IT) strategy. By examining the results in Tables 10 and 11, we can discern the efficacy of the proposed methodologies in safeguarding privacy.